Graph each circle. label the center and radius lenght. use a grafh paper
(x-3)^2 + y^2= 36

Found 2 solutions by lwsshak3, solver91311:
Answer by lwsshak3(11628)

Graph each circle. label the center and radius lenght. use a grafh paper
(x-3)^2 + y^2= 36
**
This is an equation of a circle of standard form: (x-h)^2+(y-k)^2=r^2, (h,k) being the (x,y) coordinates of the center, r=radius
For given equation:
center: (3,0)
radius=√36=6
I don't have the means to draw a circle on graph paper, but the graph below should give you a good idea how to do it.
..
y=±(36-(x-3)^2)^.5

Answer by solver91311(24713)





is a circle centered at with radius

Hint, re-write your equation:




John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism